# When Robotics Meets Wireless Communications: An Introductory Tutorial

### v3

@@ II-A1 Forward Kinematic Model
Note that while the model (2) is not invertible, the model (3) is invertible. This implies that we can fully set independently the angular speed and translational speed of the robot, but we cannot fully set indpendently its angular speed and its translational velocity.
@@ III-B UAV Channel Models
The channel models described in the previous section were originally developed for mobile communications considering ground users and are well-suited to ground MRs applications. In general, they are not appropriate for UAVs applications mainly because of the changing altitude of the UAV. Given the increasing attention being paid to the integration of UAVs into 5G networks [18, 19], for the sake of completeness of this tutorial, we next discuss briefly the modeling of UAV wireless channels, which is currently an active research topic [95, 96]. Before presenting the UAV channel models, let us discuss some particularities of UAV communications. For small-sized UAVs, the receiver is close to the UAV’s power electronics and to the motors which run continuously. As a consequence, sometimes the UAV’s motors generate electromagnetic noise that interferes with the UAV’s own receiver [97].
@@ III-B1 Air-to-ground channels
The model (116)-(117) does not take into account small-scale fading. As mentioned in [96], small-scale fading in air-to-ground channels often follows a Ricean distribution whose properties depend on the UAV altitude and the surroundings of the ground node. There are also approaches which use 3D numerical maps of the region of operation to determine the channel model [105]. We refer the reader to the survey in [96] for more detailed information about the air-to-ground UAV communications channels.
@@ IV Communications-aware Trajectory Planning
From the application point of view, a CaTP problem can be cast as either a CaR or a RaC task. We can also classify CaTP problems depending on what is being optimized into three categories. In the first category, the objective is to design the MR’s full trajectory, i.e., path and velocity profile; see for example [4]. In the second category, the translational velocity is given (e.g. constant speed) and the objective is to design the MR’s path; see for example [113]; this is a CaPP problem. In the third category, the path is given and the objective is to design the translational speed profile with which the MR will follow the given path; see for example [108].
@@ V CaTP Examples
The optimization problem is cast as a maximization problem with the following cost functional:
@@
where $\mathbf{r}_{k}\in\mathbb{R}^{3}$ is the position of user $k$, $\zeta$ captures the antenna gain, $h_{k}^{i}[n]\in\mathbb{C}$ is the channel gain between the UAV and user $k$ on subcarrier $i$ and it accounts for the multipath fading and the shadowing, $n_{k}^{i}[n]\in\mathbb{C}$ is the noise generated at user $k$’s receiver on subcarrier $i$, and $d_{k}^{i}[n]\in\mathbb{C}$ denotes the symbol transmitted from the UAV to user $k$ through subcarrier $i$th. Since $d_{k}^{i}[n]$ and $n_{k}^{i}[n]$ are stochastic processes, (168) would be an incomplete channel model without including the their statistical distributions. In [29], the authors assume that $\mathbb{E}[d_{k}^{i}[n]]=1$ and that $n_{k}^{i}[n]$ is a zero-mean Additive White Gaussian Noise (AWGN) with variance $\sigma^{2}$. They describe an offline trajectory planning problem and its online version. Here, we discuss the offline version only, where a non-causal knowledge of $h_{k}^{i}[n]$ is assumed [^15] In the online version of the problem, which is a more practical, only causal knowledge can be leveraged to design the trajectory, thus making this term deterministic; see [29] for the details of this issue.
@@
The acceleration, the speeds, and the altitude of the UAV are constrained as follows:
@@ V-C Other Trajectory Design Strategies
We now discuss some other types of CaTP problems that represent interesting and exciting research opportunities. Additional examples of CaTP can be found in [171].
@@
V-D Useful numerical tools to solve CaTP problems
@@
Finally, when the optimization target of the CaTP problem is a functional and the optimization problem involves optimizing a function instead of a set of variables, then CasADi [201], an open-source tool for nonlinear optimization and algorithmic differentiation, could be helpful. CasADi is a tool commonly used in optimal control problems. Alternatively, reinforcement learning, applicable to control problems [202], can be implemented using dedicated Python libraries.
@@ VII Discussion and Concluding Thoughts
[Uncaptioned image] Mounir Ghogho (F’18) received the M.Sc. degree in 1993 and the PhD degree in 1997 from the National Polytechnic Institute of Toulouse, France. He was an EPSRC Research Fellow with the University of Strathclyde (Scotland), from Sept 1997 to Nov 2001. In Dec 2001, he joined the school of Electronic and Electrical Engineering at the University of Leeds (England), where he was promoted to full Professor in 2008. While still affiliated with the University of Leeds, in 2010 he joined the International University of Rabat (Morocco) where he is currently Dean of Doctoral College and Director of TICLab (ICT Research Laboratory). He was awarded the UK Royal Academy of Engineering Research Fellowship in 2000 and the IBM Faculty Award in 2013. He was elevated to the grade of IEEE Fellow in 2018. His research interests are in signal processing, machine learning, and wireless communication. In the past, he served as an associate editor of the Signal Processing Magazine, the IEEE Transactions on Signal Processing, the IEEE Signal Processing Letters, and the Elsevier Digital Signal Processing journal, and a member of the IEEE Signal Processing Society SPCOM Technical Committee, the IEEE Signal Processing Society SPTM Technical Committee, and the IEEE Signal Processing Society SAM Technical Committee. He is currently a member of the steering committee of the IEEE Transactions of SIgnal and Information Processing over Networks. He held invited scientist/professor positions at Telecom Paris-Tech (France), NII (Japan), BUPT (China), University Carlos 3rd of Madrid (Spain), ENSICA (Toulouse), Darmstadt Technical University (Germany), and Minnesota University (USA).
[Uncaptioned image] Martin Saska received his MSc. degree at Czech Technical University in Prague, 2005, and his Ph.D. degree at University of Wuerzburg, Germany, within the PhD program of Elite Network of Bavaria, 2009. Since 2009, he is a research fellow at Czech Technical University in Prague, where he founded and heads the Multi-robot Systems lab (http://mrs.felk.cvut.cz/) and co-founded the Center for Robotics and Autonomous Systems with more than 70 researchers cooperating in robotics (https://robotics.fel.cvut.cz/cras/). He was a visiting scholar at the University of Illinois at Urbana-Champaign, USA in 2008, and at University of Pennsylvania, USA in 2012, 2014, and 2016, where he worked with Vijay Kumar’s group within GRASP lab. He is an author or co-author of >150 publications in peer-reviewed conferences with multiple best paper awards and more >40 publications in impacted journals, including IJRR, AURO, JFR, ASC, EJC, with >4000 citations indexed by Scholar and H-index 35. His team won multiple robotic challenges in MBZIRC 2017, MBZIRC 2020, and DARPA SubT competitions (http://mrs.felk.cvut.cz/projects/mbzirc, http://mrs.felk.cvut.cz/mbzirc2020, http://mrs.felk.cvut.cz/projects/darpa).

### v5

@@ II-A1 Forward Kinematic Model
Note that while the model (2) is not invertible, the model (3) is invertible. This implies that we can fully set independently the angular speed and translational speed of the robot, but we cannot fully set independently its angular speed and its translational velocity.
@@ III-B UAV Channel Models
The channel models described in the previous section were originally developed for mobile communications considering ground users and are well-suited to ground MRs applications. In general, they are not appropriate for UAVs applications mainly because of the changing altitude of the UAV. Given the increasing attention being paid to the integration of UAVs into 5G networks [18, 19], for the sake of completeness of this tutorial, we next discuss briefly the modeling of UAV wireless channels, which is currently an active research topic [95, 96]. Before presenting the UAV channel models, let us discuss some particularities of UAV communications. For small-sized UAVs, the receiver is close to the UAV’s power electronics and to the motors which run continuously. As a consequence, sometimes the UAV’s motors generate electromagnetic noise that can interfere with the UAV’s own receiver [97].
@@ III-B1 Air-to-ground channels
The model (116)-(117) does not take into account small-scale fading. As mentioned in [96], small-scale fading in air-to-ground channels often follows a Rician distribution whose properties depend on the UAV altitude and the surroundings of the ground node. There are also approaches which use 3D numerical maps of the region of operation to determine the channel model [105]. We refer the reader to the survey in [96] for more detailed information about the air-to-ground UAV communications channels.
@@ III-E Final comments on communications models
We conclude this section by bringing to the attention of the reader that one of the most determinant elements for the adequate selection of the communications channel model is the environment where the system will operate. To illustrate this, lets consider a pair of UAVs hovering at low altitude over a body of water while communicating. On one hand, if the body of water is a calm lake then it will act as a mirror and reflect well the electromagnetic waves. Consequently, the two-ray model described in section III-A1 could be a good selection for the communications channel model since it would consider the specular reflection occurring on the surface of the lake. On the other hand, if the body of water is a rough sea, then the irregular and dynamic surface of the rough sea would provoke that the characteristics and the number of reflected waves would also be time variant. But the two-ray model would fail to consider this, and thus it would most probably be highly inaccurate in this case. Hence the importance of taking into account the environment for the selection of the communications channel model.
@@ IV Communications-aware Trajectory Planning
From the application point of view, a CaTP problem can be categorized as either a CaR or a RaC task. This classification pertains only the nature of the application and it is not related to the mathematical formulation directly. In [113], the authors present the similar categorization in the context of UAV communications, although they don’t use the terms CaR and RaC, and they describe both categories very concisely: ’what can UAVs do for networks’ and ’what can networks do for UAVs. Thus, we can follow a similar strategy, and say that RaC are applications that result from the question ’what can robots do for networks?’ while CaR are applications that result from the question ’what can networks do for robots?’.
We can also classify CaTP problems depending on what is being optimized into three categories. In the first category, the objective is to design the MR’s full trajectory, i.e., path and velocity profile; see for example [4]. In the second category, the translational velocity is given (e.g. constant speed) and the objective is to design the MR’s path; see for example [114]; this is a CaPP problem. In the third category, the path is given and the objective is to design the translational speed profile with which the MR will follow the given path; see for example [108].
@@ V CaTP Examples
The optimization problem is cast as a minimization problem with the following cost functional:
@@
where $\mathbf{r}_{k}\in\mathbb{R}^{3}$ is the position of user $k$, $\zeta$ captures the antenna gain, $h_{k}^{i}[n]\in\mathbb{C}$ is the channel gain between the UAV and user $k$ on subcarrier $i$ and it accounts for the multipath fading and the shadowing, $n_{k}^{i}[n]\in\mathbb{C}$ is the noise generated at user $k$’s receiver on subcarrier $i$, and $d_{k}^{i}[n]\in\mathbb{C}$ denotes the symbol transmitted from the UAV to user $k$ through subcarrier $i$th. Since $d_{k}^{i}[n]$ and $n_{k}^{i}[n]$ are stochastic processes, (168) would be an incomplete channel model without including their statistical distributions. In [29], the authors assume that $\mathbb{E}[d_{k}^{i}[n]]=1$ and that $n_{k}^{i}[n]$ is a zero-mean Additive White Gaussian Noise (AWGN) with variance $\sigma^{2}$. They describe an offline trajectory planning problem and its online version. Here, we discuss the offline version only, where a non-causal knowledge of $h_{k}^{i}[n]$ is assumed [^15] In the online version of the problem, which is more practical, only causal knowledge can be leveraged to design the trajectory., thus making this term deterministic; see [29] for the details of this issue.
@@
The acceleration, the speed, and the altitude of the UAV are constrained as follows:
@@ V-C Other Trajectory Design Strategies
We now discuss some other types of interesting and exciting CaTP problems, and in the next subsection we briefly present some research opportunities in this field. Additional examples of CaTP can be found in [172].
@@
V-D Research opportunities
Some of the key technologies for UAV communications in 6G are [^16] See [113] and the references therein for a more detailed discussion on these technologies in the context of UAV communications. mmWave UAV communications, THz UAV communications, and UAV communications with Reconfigurable Intelligent Surfaces (RIS). These technologies create new research opportunities in CaTP for UAVs.
mmWave UAV communications uses highly directional antennas. This high directionality makes those communications links very sensitive to antenna misalignments. Thus, the tilting that underactuated UAVs experience when moving, see Fig. 4, creates a challenge that must be overcome if we plan to use mmWave communications in flying UAVs. This requires the development of new trajectory planning techniques, adapted to the large bandwidth of mmWave communications and focused in the antenna misalignment caused by the UAV tilting, similar to the one studied in [98].
THz communications uses extremely small wavelengths and large bandwidths. These extremely small wavelengths allow to equip small UAVs with antenna arrays. This calls for the development of new techniques for the joint optimization of trajectory planning and beamforming for UAVs equipped with antenna arrays.
RIS are passive planar structures electronically manipulated that control how they reflect the electromagnetic waves. According to [113], there is still a lack of trajectory planning techniques for UAVs equipped with RIS to improve the communications network.
Omnidirectional UAVs [197] is another interesting technology that opens new research opportunities in CaTP. These UAVs can control its position and orientation independently, as opposed to underactuated multirotor UAVs which are the focus of most of the current research in CaTP nowadays. This new capability introduced by omnidirectional UAVs can be exploited in creative ways to develop new techniques such as making a UAV orient itself in the air to direct the nul of its antenna radiation pattern towards a jammer to neutralize its attack, see [190].
Finally, all the CaTP described in the previous sections can be revisited using these new technologies while considering both robotics and communications aspects to create new and exciting developments.
V-E Useful numerical tools to solve CaTP problems
@@
Finally, when the optimization target of the CaTP problem is a functional and the optimization problem involves optimizing a function instead of a set of variables, then CasADi [203], an open-source tool for nonlinear optimization and algorithmic differentiation, could be helpful. CasADi is a tool commonly used in optimal control problems. Alternatively, reinforcement learning, applicable to control problems [204], can be implemented using dedicated Python libraries such as TensorFlow or PyTorch.
@@ VII Discussion and Concluding Thoughts
[Uncaptioned image] Mounir Ghogho Mounir Ghogho (F’18) has received the PhD degree in 1997 from the National Polytechnic Institute of Toulouse (France). He was an EPSRC Research Fellow with the University of Strathclyde (Scotland) from Sept 1997 to Nov 2001. In Dec 2001, he joined the University of Leeds (England), where he was promoted to full Professor in 2008. While still affiliated with the University of Leeds, in 2010 he joined the International University of Rabat (UIR) where he is currently Dean of the College of Doctoral Studies and Director of TICLab (ICT Research Lab). He is a Fellow of IEEE and AAIA (Asia-Pacific AI Association), a recipient of the 2013 IBM Faculty Award, and a recipient of the 2000 UK Royal Academy of Engineering Research Fellowship. His research interests are in Machine Learning, Signal Processing and Wireless Communication. He has coordinated around 20 research projects and supervised over 50 PhD students in the UK and Morocco. He has served as Associate Editor of many journals including the IEEE Signal Processing Magazine and the IEEE Transactions on Signal Processing. He has also been a member of The Signal Processing Theory and Methods (SPTM) and The Sensor Array and Multichannel (SAM) technical committees of the IEEE Signal Processing Society. He has chaired many international conferences and workshops including the IEEE 11th Workshop on Signal Processing for Advanced Wireless Communication (SPAWC 2010) and the 21st European Signal Processing Conference (EUSIPCO 2013).
[Uncaptioned image] Martin Saska received his Ph.D. degree at University of Wuerzburg, Germany, within PhD program of Elite Network of Bavaria. He founded and heads the Multi-robot Systems lab (http://mrs.felk.cvut.cz/) at Czech Technical University in Prague with more than 40 researchers. He was a visiting scholar at University of Illinois at Urbana-Champaign and at University of Pennsylvania, USA. He is co-author of >200 publications in conferences and impacted journals, including IJRR, AURO, JFR, ASC, EJC, with >6800 citations indexed by Scholar and H-index 45. His team won multiple robotic challenges in MBZIRC 2017, MBZIRC 2020 and DARPA SubT competitions (http://mrs.felk.cvut.cz/projects/mbzirc, http://mrs.felk.cvut.cz/mbzirc2020, http://mrs.felk.cvut.cz/projects/darpa).